OpenStudy (inowalst):

Which of the following is the rational exponent expression of fifth root of 7 n?

OpenStudy (inowalst):

That's the one in the question:\[\sqrt[5]{7n}\]

OpenStudy (inowalst):

\[5n ^{7}\]\[(7n)^{5}\]\[7n ^{\frac{ 1 }{ 5 }}\]\[(7n)^{\frac{ 1 }{ 5 }}\]

OpenStudy (inowalst):

I chose the last one.

OpenStudy (anonymous):

The last one

OpenStudy (anonymous):

\[\large \color{green}{\sqrt[n]{m} = (m)^{\frac{1}{n}}}\]

OpenStudy (inowalst):

Okay, thank you.

OpenStudy (theraggedydoctor):

dot.